OpenStudy (anonymous):

If 10,800 cm2 of material is available to make a box with a square base and an open top, find the largest possible volume of the box.

OpenStudy (anonymous):

now this you need calc for

OpenStudy (anonymous):

put base at x, height as h and so you know that \[x^2+4xh=10800\] \[h=\frac{10800-x^2}{4x}\] \[V(x)=x^2h=x^2(\frac{10800-x^2}{4x})\] \[V(x)=x\frac{10800-x^2}{4}=2700x-\frac{x^3}{4}\]

OpenStudy (anonymous):

now to find the max take the derivative, set = 0 and solve, get \[V'(x)=2700-\frac{3}{4}x^2\] then \[2700-\frac{3}{4}x^2=0\] \[\frac{3}{4}x^2=2700\] \[x^2=3600\] \[x=60\]

OpenStudy (anonymous):

how'd i do?

OpenStudy (anonymous):

\[\color{red}{\text{any typos my myininaya?}}\]

OpenStudy (anonymous):

hmmm i dont believe thats correct, satellite73. My teacher says 60 is not the answer :/